OpenStudy (calculusxy):

Tom had 350 dollars to spend on 9 books. After buying them he had 17 dollars. How much did each book cost. Write an equation, then evaluate it.

OpenStudy (anonymous):

350-17= 333, 333/9 = your answer

OpenStudy (calculusxy):

You mean by $37?

OpenStudy (anonymous):

?

OpenStudy (calculusxy):

Is the answer $37? I got this answer before, but wanted to check.

OpenStudy (lena772):

350-17=money spent on books (333) 333/9= cost of each book(33)

OpenStudy (calculusxy):

Another one guys and gals?

OpenStudy (lena772):

cost of each book (37) **

OpenStudy (lena772):

Post & tag me

OpenStudy (anonymous):

thank you @Lena772 , and sure

OpenStudy (calculusxy):

Yup

OpenStudy (calculusxy):

@Mel98 I tagged you to my new question. I need help!

OpenStudy (lena772):

Tag me if you want

OpenStudy (calculusxy):

@I tagged you